# System-Theoretic Analysis of Dynamic Generalized Nash Equilibrium Problems – Turnpikes and Dissipativity

### v1

@@ Abstract
Generalized Nash equilibria are used in multi-agent control applications to model strategic interactions between agents that are coupled in the cost, dynamics, and constraints. We study the properties of open-loop GNE trajectories from a system-theoretic perspective. We show how strict dissipativity generates the turnpike phenomenon in GNE solutions. Moreover, we establish a converse turnpike result, i.e., the implication from turnpike to strict dissipativity. We derive conditions under which the steady-state GNE is the optimal operating point and, using a game value function, we give a local characterization of the geometry of storage functions. Finally, we design linear terminal penalties that ensure GNE open-loop trajectories converge to and remain at the steady-state GNE. These connections provide the foundation for future system-theoretic analysis of GNEs similar to those existing in optimal control.
@@
dynamic games, optimal control, Nash equilibrium problems, multi-agent systems, dissipativity, turnpike property, non-cooperative game theory
@@ 1 Introduction
Control is at the core of many multi-agent applications managing interactions between strategic agents with conflicting objectives and coupled action spaces. Generalized Nash equilibria (GNEs) have emerged as a promising solution concept at the intersection of game theory and optimal control [Facchinei and Kanzow, 2009]. They naturally enable to define per-agent optimal control problems that are coupled, in a game-theoretic sense, in the objective, constraints, and dynamics. GNEs have been applied to resource allocation problems (e.g., energy [Atzeni et al., 2012, Hall et al., 2022], transportation [Bassanini et al., 2002], telecommunications [Pavel, 2012]) and competitive settings (e.g., autonomous driving [Dreves and Gerdts, 2017, Le Cleac’h et al., 2022], supply chains [Hall et al., 2024a]).
@@
The prospect for applications has catalyzed extensive research on efficient GNE-seeking algorithms and their convergence guarantees [Facchinei et al., 2010, Belgioioso et al., 2022]. The resulting solution sets have been studied in detail [Kulkarni and Shanbhag, 2012, Nabetani et al., 2009], as well as their efficiency [Kulkarni, 2019] and fairness properties [Hall et al., 2025].
@@
A fundamental step towards safe deployment of controllers is the characterization of their closed-loop properties. Yet, the behavior of finite-horizon trajectories resulting from GNEs remains poorly understood, even in the open-loop case. Previous works have applied dissipativity techniques to analyze the feedback mapping [Hall et al., 2024b] and drawn connections to dynamic variational inequalities [Benenati and Grammatico, 2024]. Despite this progress finite-horizon, GNE trajectories lack the rich system-theoretic characterizations developed for optimal control problems over the past decades, cf. [Willems, 1971, Wilde and Kokotovic, 1972, Trélat and Zuazua, 2015] and many others.
In classic optimal control, necessary first-order conditions are given by the maximum principle [Boltyanskii et al., 1960] while dynamic programming provides necessary and sufficient conditions [Bellman, 1954]. Moreover, the turnpike phenomenon is a similarity property of parametric optimal control problems with respect to varying initial conditions and horizon lengths. In essence, the phenomenon implies that for different initial conditions and horizon lengths the optimal trajectories approach the neighborhood of an optimal operating point (entry arc), remaining there (turnpike arc) and then diverge from it towards the end of the horizon (leaving arc). First observations date back to [Ramsey, 1928] and [von Neumann, 1938], while the name turnpike was coined by [Dorfman et al., 1958]. We refer to [McKenzie, 1976] and [Faulwasser and Grüne, 2022] for classic and modern literature overviews. A recent key development is the analysis of the links of the turnpike phenomenon with dissipativity properties [Grüne and Müller, 2016, Faulwasser et al., 2017], which is in turn linked to the dissipativity route to infinite-horizon optimal control which originated with [Willems, 1971], see [Faulwasser and Kellett, 2021] for recent nonlinear extensions. For results on turnpike analysis exploiting the structures of the maximum principle we refer to [Trélat and Zuazua, 2015]. In applications such as energy management, the turnpike effect or end-of-horizon effects lead, for instance, to myopic discharging of batteries [Hall et al., 2022, Vincent et al., 2020].
@@
Interestingly, turnpike phenomena have also been observed in games in the Economics literature as early as the 1980s [Fershtman and Muller, 1986] and was later treated from a theoretical perspective for infinite-horizon open-loop games in continuous time [Carlson and Haurie, 1995] and discrete time [Carlson and Haurie, 1996] focusing on existence, uniqueness of equilibria and convergence to the turnpike. Extensions to differential games with coupled state constraints followed [Carlson and Haurie, 2000].
The turnpike property in games has recently received renewed attention. Turnpike properties for stochastic differential LQ games are studied in [Li et al., 2025] for the zero-sum setting and in [Cohen and Jian, 2025] for the non-zero sum setting. Multiple works derive turnpike properties for mean field games under a large population assumption [Cirant and Porretta, 2021, Carmona and Zeng, 2024, Ersland et al., 2025, Fedorov, 2025]. Further, turnpikes were observed in competitive supply chain games [Hall et al., 2024a].
@@
This paper builds a bridge between the dissipativity route to turnpikes in optimal control and dynamic games. We establish the turnpike property for discrete-time finite-horizon GNE problems and study its implications on dissipativity of the GNE with respect to the steady-state GNE. Specifically, we make the following four contributions:
@@
    Demonstrating the crucial structural link between turnpike properties in parametric OCPs and parametric GNEs providing a foundation for system-theoretic analysis of GNEs.
@@
    Establishing that strict dissipativity implies the turnpike property, also in GNEs. Further, we derive a converse turnpike result, connecting the turnpike of the open-loop trajectory to dissipativity with respect to the steady-state GNE.
@@
    Providing an optimality-like interpretation of GNEs by considering a game value function and studying its local gradient structure. Additionally, we show that the storage function’s gradient equals the sum of agents’ dual multipliers at the steady-state GNE.
@@
    Designing mechanisms to suppress the leaving arc in GNE open-loop trajectories, using per-agent linear terminal penalties to ensure trajectories converge to and remain at the optimal operating point.
@@
2 Problem Setting: Finite-Horizon Generalized Nash Equilibrium Problem
A group of self-interested agents $v\in\mathcal{V}:=\{1,\dots,M\}$ controls the following shared dynamics
$x_{k+1}=f(x_{k},u_{k}^{v},u_{k}^{-v})$ (1)
for a compact set of initial conditions $\mathbf{x}\in\mathbb{X}_{0}$. Each agent minimizes their objective function over a prediction horizon $N$ with states and inputs coupled to other agents. This constitutes a finite-horizon Generalized Nash Equilibrium problem (GNEP) as follows
@@ 2 Problem Setting: Finite-Horizon Generalized Nash Equilibrium Problem
$G^{v}\,u_{k}^{v}\leq h^{v},\hskip 54.00002ptk\in\mathbb{Z}_{N}$ (2d)
$\;x_{0}=\mathbf{x}.$ (2e)
where $\mathbf{x}$ is the initial condition of system (1). We define the following per-agent and global action sets
@@
and the cumulative cost
$J_{N}^{v}(x,u):=\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}).$
Decisions that jointly solve (2) are generalized Nash equilibria (GNE) [Facchinei and Pang, 2009, §2], as defined next.
@@
A joint decision $(x^{*},u^{*})$ is a GNE of (2) if $\forall v\in\mathcal{V}$
$J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})$
$\forall(x,u^{v})\in\mathcal{Z}(\mathbf{x},u^{-v*})$. The corresponding solution set for fixed $N\in\mathbb{N}$ and $\mathbf{x}\in\mathbb{X}_{0}$ is denoted as
@@
Intuitively, a GNE is a joint decision for which no agent $v\in\mathcal{V}$ can reduce its cost by unilaterally changing its own decision.
For the sake of concise exposition and in analogy to optimal control concepts, we refer to $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ as a GNE pair or an optimal game pair. Moreover, with slight abuse of notation we write
$(x_{k},u_{k})\in\Pi_{k}\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$
to express that there exist $k\in\mathbb{Z}_{N}$ and $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ such that
$(x_{k},u_{k})=(x_{k}^{*},u_{k}^{*})\quad\text{with}\quad(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}).$
We want to study the convergence of the finite-horizon GNE trajectory with respect to a steady-state point. Thus, we introduce the game-theoretic equilibrium of the steady-state game next.
@@
The pair $(x_{s},u_{s})$ is called a steady-state GNE if it solves the steady-state GNEP
@@
with the corresponding solution set $\mathcal{S}^{\text{\tiny GNE}}_{s}\subset\mathbb{R}^{n_{x}+n_{u}}.$
We analyze properties of the game pair of the GNE in (2) provided a solution exists. Thus, in the following we assume that $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and $\mathcal{S}^{\text{\tiny GNE}}_{s}$ are nonempty $\forall\mathbf{x}\in\mathbb{X}_{0}$.
Remark 1 (Existence & computation).
The theoretical results which follow hold for the nonlinear GNE problem in (2). Existence of a GNE has been established under a variety of assumptions as summarized in [Dutang, 2013]. Efficient algorithms to compute solutions of nonlinear GNEPs are a topic of ongoing research, one example is given in [Le Cleac’h et al., 2022].
3 Dissipativity and Turnpike Relations for GNEPs
Turnpike and dissipativity properties are closely linked, providing system-theoretic insights for optimal control. In the following, we demonstrate that this link also exists for the GNE problem in (2). We first introduce a performance measure for the population of agents as follows
@@ 3 Dissipativity and Turnpike Relations for GNEPs
and
$J_{N}(x,u):=\sum_{k=0}^{N-1}\ell(x_{k},u_{k})=\sum_{k=0}^{N-1}\sum_{v\in\mathcal{V}}\ell^{v}(x_{k},u_{k}^{v},u_{k}^{-v}).$
While using the sum of costs may seem unusual from a game-theoretic perspective, [Benenati and Grammatico, 2024] show that it can serve as a Lyapunov function for LQ games with shared constraints under suitable controllability assumptions. This motivates our choice in (11) and simplifies the subsequent theoretical analysis.
3.1 Dissipativity of GNEPs
Based on the joint performance measure, we can make a strict dissipativity assumption analogous to the ones existing in optimal control.
@@ 3.1 Dissipativity of GNEPs
Given a steady-state GNE $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$, the GNEP (2) is called strictly dissipative if for the supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ there exists a storage function $\Lambda:\mathbb{X}_{0}\to\mathbb{R}$ bounded from below and $\alpha_{\ell}\in\mathcal{K}$ such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$
@@
holds for each $(x_{k},u_{k})\in\Pi_{k}\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$.
Figure 1: Schematic of a state trajectory exhibiting the turnpike property for different horizon lengths.
[Willems, 1972] introduced equivalent dissipativity characterizations via the required supply and the available storage. In discrete time the available storage characterization can be traced back to [Byrnes and Lin, 1994]. For our further investigations, the latter one is very helpful. In terms of GNEPs, this well-known characterization reads as follows.
Theorem 1 (Available storage of GNEP ).
The GNEP (2) is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ and $\alpha_{\ell}\in\mathcal{K}_{\infty}$ if and only if
$\Lambda_{a}(\mathbf{x})=$ $\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x^{*},u^{*})\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}\sum_{k=0}^{N-1}-s(x_{k}^{*},u_{k}^{*})+\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)<\infty.$
Note that Definition 3 and Theorem 1 consider strict dissipativity only along game pairs resulting from the GNEP in (2). In optimal control, strict dissipativity is often assumed to hold along all feasible trajectories, see for instance [Grüne and Müller, 2016, Defn. 2.1], while [Faulwasser et al., 2017] also restricted the crucial inequality to hold only along optimal trajectories. However, in a game-theoretic setting we need to restrict the definition as the following shows.
Remark 2 (Why restrict dissipativity to $\mathcal{S}^{\text{\tiny GNE}}_{N}?$).
Denote the optimal steady state associated with (11) as

### v3

# System-Theoretic Analysis of Dynamic Generalized Nash Equilibria – Turnpikes and Dissipativity

@@ Abstract
Generalized Nash equilibria are used in multi-agent control applications to model strategic interactions between agents that are coupled in the cost, dynamics, and constraints, and provide the foundations for game-theoretic MPC (Receding Horizon Games). We study properties of finite-horizon dynamic GNE trajectories from a system-theoretic perspective. We show how strict dissipativity generates the turnpike phenomenon in GNE solutions. Moreover, we establish a converse turnpike result, i.e., the implication from turnpike to strict dissipativity. We derive conditions under which the steady-state GNE is the optimal operating point and, using a game value function, we give a local characterization of the geometry of storage functions. Finally, we design linear terminal penalties that ensure dynamic GNE trajectories applied in open-loop converge to and remain at the steady-state GNE. These connections provide the foundation for future system-theoretic analysis of GNEs similar to those existing in optimal control as well as for recursive feasibility and closed-loop stability results of game-theoretic MPC.
@@
dynamic games, optimal control, generalized Nash equilibrium problems, multi-agent systems, dissipativity, turnpike property, non-cooperative game theory
@@ 1 Introduction
Control is at the core of many multi-agent applications managing interactions between strategic agents with conflicting objectives, dynamics, and coupled action spaces. Generalized Nash equilibria (GNEs) have emerged as a promising solution concept at the intersection of game theory and optimal control [24]. They naturally enable to define per-agent optimal control problems that are coupled, in a game-theoretic sense, in the objective, constraints, and dynamics. GNEs have been applied to resource allocation problems in dynamic settings (e.g., energy [4, 36], transportation [5], telecommunications [51]) and competitive dynamic settings (e.g., autonomous driving [21, 43], supply chains [38]).
@@
Applications in real-time control of dynamic multi-agent systems catalyzed extensive research on efficient GNE-seeking algorithms and their convergence guarantees [32, 6]. The resulting solution sets have been studied in detail [41, 48], as well as their efficiency [42] and fairness properties [37].
@@
A fundamental step towards safe deployment of controllers is the characterization of their closed-loop properties. Yet, the behavior of finite-horizon trajectories resulting from GNEs remains poorly understood, even in the open-loop case. Previous works have applied dissipativity techniques to analyze the feedback mapping [35] and drawn connections to dynamic variational inequalities [8]. Despite this progress, finite-horizon GNE trajectories lack the rich system-theoretic characterizations developed for optimal control problems over the past decades, cf. [57, 2, 54] and many others.
Figure 1: Schematic of a dynamic GNE state trajectory $x^{*}$ exhibiting the turnpike property at the steady-state GNE $x_{s}$ for different horizon lengths.
In this context, the turnpike phenomenon arises as a property of solutions to an optimal control problem (OCP) solved for different initial conditions and varying horizon lengths. The phenomenon is characterized by the solutions clustering near one common steady state which is referred to as the turnpike, see Figure 1. It is a similarity property of parametric OCPs and thus is a crucial element in the closed-loop analysis of OCPs applied in a receding-horizon fashion, i.e., in model predictive control (MPC). First observations date back to [52] and [56], while the notion turnpike was coined by [20]. We refer to [45] and [28] for classic and modern literature overviews. A recent key development is the uncovering of the close connection between the turnpike phenomenon and dissipativity properties [33, 27], which is in turn linked to the dissipativity route to infinite-horizon optimal control which originated with [57], see [26] for recent nonlinear extensions. The link between turnpike and dissipativity enabled major breakthroughs in the closed-loop stability analysis of nonlinear and economic MPC, see, e.g., [34]. In applications such as energy management, the turnpike effect or end-of-horizon effect leads, for instance, to myopic discharging of batteries [36, 55].
@@
Interestingly, turnpike phenomena have also been observed in games in the Economics literature as early as the 1980s [31] and were later treated from a theoretical perspective for infinite-horizon open-loop games in continuous time [11] and discrete time [12] focusing on existence, uniqueness of equilibria and convergence to the turnpike. Extensions to differential games with coupled state constraints followed [13]. Turnpike properties in games have recently received renewed attention: The specialized class of stochastic differential LQ games with continuous-time dynamics are studied in [44] for the zero-sum setting and in [17] for the non-zero sum setting. Multiple works derive turnpike properties for mean field games under a large population assumption [16, 14, 23, 30]. Further, turnpikes were observed in competitive dynamic supply chains [38].
@@
This paper lays the foundations for an overarching theory of turnpike and dissipativity analysis in noncooperative games for a general class of cost functions, constraints, and discrete-time dynamics. We build a bridge between the dissipativity route to turnpikes in optimal control and dynamic games. Specifically, we establish the turnpike property for finite-horizon discrete-time dynamic GNE problems and study its implications on dissipativity of the GNE with respect to the steady-state GNE. Overall, we make the following four contributions:
@@
    We demonstrate the crucial structural link between turnpike properties in parametric OCPs and parametric GNEs providing a foundation for system-theoretic analysis of GNEs.
@@
    We establish that strict dissipativity implies the turnpike property, also in GNEs. Further, we derive a converse turnpike result, connecting the turnpike of the open-loop trajectory to dissipativity with respect to the steady-state GNE.
@@
    We provide an optimality-like interpretation of GNEs by considering a game value function and studying its local gradient structure. Additionally, we show that the storage function’s gradient equals the sum of agents’ dual multipliers at the steady-state GNE.
@@
    We design mechanisms to suppress the leaving arc in GNE open-loop trajectories, using per-agent linear terminal penalties to ensure trajectories converge to and remain at the steady-state GNE.
@@
2 Finite-Horizon Generalized Nash Equilibrium Problems
Consider a nonlinear discrete-time system
$x_{k+1}=f(x_{k},u_{k}),\quad x_{0}=\mathbf{x}$
with state $x_{k}\in\mathbb{R}^{n_{x}}$, control input $u_{k}\in\mathbb{R}^{n_{u}}$, and initial condition $\mathbf{x}$ which lies in a compact set $\mathbf{x}\in\mathbb{X}_{0}$. Suppose the input is not computed by a central entity but is instead a stacked vector $u_{k}=[(u_{k}^{1})^{\top},\ldots,(u_{k}^{M})^{\top}]^{\top}$, where each component $u_{k}^{v}$ is decided by a self-interested agent $v\in\mathcal{V}:=\{1,\dots,M\}$. Each agent aims to steer the shared dynamics in its own favor
$x_{k+1}=f(x_{k},u_{k}^{v},u_{k}^{-v}),\quad x_{0}=\mathbf{x}.$ (1)
Specifically, each agent $v$ minimizes its accumulated stage cost $\ell^{v}$ over a prediction horizon $N$ subject to the shared dynamics (1). In addition to this dynamic coupling, agents influence each other’s outcomes through coupling in their cost and constraints, while still being subject to individual local constraints. Formalizing this gives rise to a finite-horizon dynamic Generalized Nash Equilibrium Problem (GNEP) of the form
@@ 2 Finite-Horizon Generalized Nash Equilibrium Problems
$h^{v}(u_{k}^{v})\leq 0,\hskip 56.00014ptk\in\mathbb{Z}_{N}$ (2d)
$\;x_{0}=\mathbf{x},$ (2e)
where (2d) are nonlinear local constraints and (2c) nonlinear coupled constraints.
For each choice of the initial condition $\mathbf{x}$ and the horizon $N$, the GNEP (2) generates finite-horizon GNE trajectories. Hence, we view the GNEP (2) as a parametric game-theoretic decision problem. Note that—for now—we make no assumptions on the functions in (2) but will add them when our theoretical results require them.
We define the following per-agent and global action sets
@@
as well as $\mathcal{Z}_{\infty}(\mathbf{x})$ for the infinite horizon problem and the per-agent cumulative cost
$J_{N}^{v}(x,u^{v},u^{-v}):=\sum_{k=0}^{N-1}\ell^{v}(x_{k},u_{k}^{v},u^{-v}_{k}).$
With these abstractions we can write (2) as follows:
$\forall v\in\mathcal{V}:\;\left\{\begin{aligned} \;\min_{u^{v},x}\;&J^{v}_{N}(x,u^{v},u^{-v})\\ \operatorname{\ {s.}{t.}\ }&(x,u^{v})\in\mathcal{Z}_{N}^{v}(\mathbf{x},u^{-v}).\end{aligned}\right$. (5)
Decisions that jointly solve (2) are called generalized Nash equilibria (GNE) [25, §2]. Intuitively, at a GNE no agent $v\in\mathcal{V}$ can reduce its cost by unilaterally changing its own decision.
@@
A joint decision $(x^{*},u^{*})\in\mathcal{Z}_{N}(\mathbf{x})$ is a GNE of (2) if
$\forall v\in\mathcal{V}:\,J_{N}^{v}(x^{*},u^{v*},u^{-v*})\leq J_{N}^{v}(x,u^{v},u^{-v*})$
$\forall(x,u^{v})\in\mathcal{Z}_{N}(\mathbf{x},u^{-v*})$. The corresponding solution set for fixed $N\in\mathbb{N}$ and $\mathbf{x}\in\mathbb{X}_{0}$ is denoted as
@@
In analogy to optimal control concepts, we refer to $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ as a GNE pair or an optimal game pair. As a shorthand, the point-wise in time projection of $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ onto the states is written as $\mathbb{X}_{N}(\mathbb{X}_{0})\subseteq\mathbb{R}^{n_{x}}$. That is, if $\tilde{x}\in\mathbb{X}_{N}(\mathbb{X}_{0})$ then there exist at least one game pair $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x}),\mathbf{x}\in\mathbb{X}_{0}$ such that the state trajectory $x^{*}$ passes through $\tilde{x}$ at least once. Similarly $X_{\infty}(\mathbb{X}_{0})$ refers to the infinite horizon problem.
Henceforth, we study convergence or clustering of finite-horizon GNE trajectories close to a two-fold equilibrium: a point that is both (i) a steady-state of (1), i.e., $\bar{x}=f(\bar{x},\bar{u}^{v},\bar{u}^{-v})$; and (ii) a strategic (decision) equilibrium of the one-step GNEP in (2) such that no agent benefits from deviating in the next time step. This is formalized in the following steady-state GNEP.
@@
The pair $(x_{s},u_{s})$ is called a steady-state GNE if it solves
@@
with the corresponding solution set $\mathcal{S}^{\text{\tiny GNE}}_{s}\subset\mathbb{R}^{n_{x}+n_{u}}.$ Problem (11) is called a steady-state GNEP.
We analyze properties of the GNE in (2) and (11) provided a solution exists. Thus, in the following we assume that $\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$ and $\mathcal{S}^{\text{\tiny GNE}}_{s}$ are nonempty $\forall\mathbf{x}\in\mathbb{X}_{0}$.
Remark 1 (Existence and computation).
The theoretical results which follow hold for the nonlinear GNE problem in (2). Existence of a GNE has been established under a variety of assumptions, as summarized in [22]. One approach is to assume that $\forall v\in\mathcal{V},\;\forall\mathbf{x}\in\mathbb{X}_{0},\;\forall u^{-v}\in\mathcal{U}^{-v}:=\Pi_{j\in\mathcal{V}\backslash v}\mathcal{U}^{j}\$the set $\mathcal{Z}_{N}^{v}(\mathbf{x},u^{-v})$ and $\mathcal{U}^{v}$ is nonempty, compact and convex and the function $J^{v}_{N}(\cdot,\cdot,u^{-v})$ is convex in both arguments [50, Prop. 12.7]. Efficient algorithms to compute solutions of nonlinear GNEPs are a topic of ongoing research, one example is given in [43].
3 Dissipativity and Turnpikes for GNEPs
Turnpike and dissipativity properties are closely linked, providing system-theoretic insights for optimal control. In the following, we demonstrate that this link also exists for the GNE problem in (2).
Consider the social welfare as a joint performance measure for the population of agents
@@ 3 Dissipativity and Turnpikes for GNEPs
and introduce the shorthand
$J_{N}(x,u):=\sum_{k=0}^{N-1}\ell(x_{k},u_{k}).$ (13)
Remark 2 (GNEPs vs. OCPs).
At this point it is fair to ask for the precise distinction between the GNE setting considered here and an optimal control point of view. Solving (5) yields a GNE, i.e., a joint decision from which no agent $v\in\mathcal{V}$ can reduce its cost unilaterally. In contrast, the optimal control counterpart of (5) reads
$\;V^{\diamond}_{N}(\mathbf{x}):=\min_{u,x}\;J_{N}(x,u)\operatorname{\ {s.}{t.}\ }(x,u)\in\mathcal{Z}_{N}(\mathbf{x},u),$
where all control actions are chosen with the sole objective of minimizing the cost for the entire agent population (13). Whenever necessary, we use superscript $\cdot^{*}$ for GNEP quantities and $\cdot^{\diamond}$ for OCP quantities.
Take a simple unconstrained quadratic example with two agents and cost $J^{1}(u^{1},u^{2})=r^{1,1}(u^{1})^{2}+r^{1,2}u^{1}u^{2}+r^{1}_{\text{lin}}u^{1}$, and analogously for agent 2. Any equilibrium solution must satisfy
$\left[\begin{smallmatrix}\frac{\partial J^{1}(u^{1},u^{2})}{\partial u^{1}}\\ \frac{\partial J^{2}(u^{1},u^{2})}{\partial u^{2}}\end{smallmatrix}\right]=\underbrace{\left[\begin{smallmatrix}2r^{1,1}&r^{1,2}\\ r^{2,1}&2r^{2,2}\end{smallmatrix}\right]}_{R^{*}}u+\left[\begin{smallmatrix}r^{1}_{\text{lin}}\\ r^{2}_{\text{lin}}\end{smallmatrix}\right]\stackrel{{\scriptstyle!}}{{=}}0$
whereas $R^{\diamond}=\left[\begin{smallmatrix}2r^{1,1}&0.5(r^{1,2}+r^{2,1})\\ 0.5(r^{1,2}+r^{2,1})&2r^{2,2}\end{smallmatrix}\right]$ and thus $u^{*}$ and $u^{\diamond}$ diverge with diverging $r^{1,2}$ and $r^{2,1}$. It is well known that in noncooperative games where agents act selfishly, the equilibrium solution does not minimize the sum of costs, and thus differs from the OCP solution. This maximal efficiency loss is referred to as the price of anarchy (PoA); see [42] for an analysis in the context of GNEs. Formally,
$\text{PoA}(\mathbf{x}):=\dfrac{\displaystyle\sup_{(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})}J_{N}(x^{*},u^{*})}{V^{\diamond}_{N}(\mathbf{x})}\geq 1.$ (14)
Next, using the joint performance measure $J_{N}$ in (13), we introduce a strict dissipativity notion which follows concepts used in optimal control.
@@
Given a steady-state GNE $(x_{s},u_{s})\in\mathcal{S}^{\text{\tiny GNE}}_{s}$, the GNEP (2) is called strictly dissipative with supply rate
$s(x_{k},u_{k}):=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$
if there exists a storage function $\Lambda:\mathbb{X}_{N}(\mathbb{X}_{0})\to\mathbb{R}$ bounded from below such that $\forall N\in\mathbb{N},\forall\mathbf{x}\in\mathbb{X}_{0}$
@@
holds for some $\alpha_{\ell}\in\mathcal{K}$ and each point $(x_{k},u_{k})$ along game pairs $(x,u)\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$. If (14) holds for $\alpha_{\ell}\equiv 0$, the GNEP (2) is called dissipative.
Since $\alpha_{\ell}$ is of class $\mathcal{K}$, inequality (14) is strict whenever $(x_{k},u_{k})\neq(x_{s},u_{s})$, hence the term strict dissipativity is used in the literature. A natural question is how to verify strict dissipativity of GNEPs. To this end, consider
$\Lambda_{\alpha_{\ell}}(\mathbf{x}):=\displaystyle\hskip-5.69054pt\sup_{\begin{subarray}{c}N\in\mathbb{N}\\ (x^{*},u^{*})\\ \in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})\end{subarray}}\hskip-2.84526pt\sum_{k=0}^{N-1}\alpha_{\ell}\left(\left\|\begin{smallmatrix}x_{k}^{*}-x_{s}\\ u_{k}^{*}-u_{s}\end{smallmatrix}\right\|\right)-s(x_{k}^{*},u_{k}^{*}).$ (15)
Observe that, given the initial condition $\mathbf{x}$, this is a free end-time optimization problem over all game pairs $(x^{*},u^{*})\in\mathcal{S}^{\text{\tiny GNE}}_{N}(\mathbf{x})$. The objective is to compute the maximum storage that can be extracted from the system, hence $\Lambda_{\alpha_{\ell}}$ is referred to as the available storage [58, 10]. The next result extends the classical characterization of dissipativity via available storage [58] to dynamic GNEPs.
Theorem 1 (Available storage of GNEPs & (14)).
The GNEP (2) is strictly dissipative with supply rate $s(x_{k},u_{k})=\ell(x_{k},u_{k})-\ell(x_{s},u_{s})$ and $\alpha_{\ell}\in\mathcal{K}$ (Definition 3) if and only if, for all $\mathbf{x}\in\mathbb{X}_{\infty}(\mathbb{X}_{0})$, the bound $\Lambda_{\alpha_{\ell}}(\mathbf{x})<\infty$ holds.
Proof 3.1.
The proof is given in Appendix A.
Willems’ motivation for the introduction of dissipativity notions in systems and control has been the “generalization of Lyapunov functions to open systems, to systems with inputs and outputs” [59]. From this perspective, the storage function is the amount of energy in an open system and the available storage the amount that can be extracted via inputs and outputs. Thus, Theorem 1 states that strict dissipativity, which loosely implies that solutions are attracted to $(x^{s},u^{s})$, can only hold if for all initial conditions the energy extractable along GNE trajectories is bounded; if it were unbounded, no such attraction could occur.
In OCPs, strict dissipativity is often assumed to hold along all feasible trajectories [33, Defn. 2.1], though [27] restrict it to optimal ones. As the next result shows, in the game-theoretic setting the restriction to GNE trajectories of (2) is indeed necessary.
Lemma 2.
(Restriction of domain of (14)) The strict dissipativity inequality (14) does not necessarily hold along all feasible trajectories $(x,u)\in\mathcal{Z}_{N}(\mathbf{x})$.
Proof 3.2.
We introduce the optimal control counterpart of the steady-state game in (11) which gives the steady-state minimizer